Main INDEX
Monthly INDEX
PREV
NEXT
User name dhamilton
Log entry time 00:05:05 on April 7,2008
Entry number 230458
keyword=Swing Shift Summary
Swing Shift Summary.
D. Hamilton/R. Lindgren (SL)
K. Allada (TO)
K. Fissum (3rd)
pi0 production kinematics.
BigBite = 54 degrees, HRSl = 20.5 degrees.
16:00 - Shift started during run 2449.
Run 2449 (0.45M)
Run 2450 (0.45M)
Run 2451 (0.45M)
Run 2452 (0.45M)
Run 2453 (0.45M)
Run 2454 - junk run due to ROC5 problems.
Run 2455 (0.45M)
Run 2456 (0.45M)
18:00 - problems with the BB HV due to problems with the beam position.
Run 2457 (0.31M) - ended early due to BB HV trip.
Run 2458 (1M) - first run with CODA file splitting enabled.
Run 2459 (1M)
Run 2460 (0.93M) - ended early due to loss of beam
19:40 - MCC called to say there were some problems with beam instability. Beam was away for 15 minutes. Several BB HV trips related to beam instability.
Run 2461 (0.66M) - ended early due to loss of beam.
Run 2462 (0.31M) - ended early due to loss of beam.
20:55 - problems persist with beam instability. MCC are trying to fix the problem.
21:10 - more HV trips when beam was restored to hall. Beam is stable. The last plane on the MWDC is drawing too much current. Suspect a broken wire.
21:30 - hall status changed to controlled access in order to identify the problem with BB MWDC back.
Shift ends with people in the hall trying to fix the problem with BB MWDC. HRSl is changing back to 960 MeV/c to continue production running. BigBite coil current is set to 0A for now.